A rugby player passes the ball 7.25 m across the field, where it is caught at the same height as it left his hand. At what angle was the ball thrown if its initial speed was 12.0 m/s, assuming that the smaller of the two possible angles was used?

angle: °

What other angle gives the same range?

other angle: °

For the pass using the smaller of the two angles, how long is the ball in the air?

length of time:
Answer #1

answer) we know the formula for the range of a projectile

R=v2sin2/g

R=122*sin2/9.8

7.25=144*sin2/9.8

sin2=9.8*7.25

sin2=71.05/144

=1/2sin-1(0.4934028)

=14.8o

so the answer is 14.8o or 15o ( this is the answer for first part)

2nd part

the other angle that gives the same range 90-14.8=75.2o

so the answer is 75.2o or 75o

the 3rd part

for time we have the formula

t=2vsin/g=2*12*sin14.8/9.8=0.625 s

so the answer is 0.625 s or 0.62 s